# A geometric approach toconjugation-invariant random permutations

### v3

@@ 1.1 Conjugation-invariant random permutations
We say that a random permutation $\tau$ of $[n]:=\{1,\dots,n\}$ is conjugation-invariant when for any fixed permutation $\rho$, the conjugate $\rho\circ\tau\circ\rho^{-1}$ follows the same law as $\tau$. Standard examples of conjugation-invariant permutations are uniformly random permutations, random involutions, Ewens random permutations [Ewe72], their generalizations [Kin78, Tsi00, BUV11], and many more.
@@
There are reasons to believe in universality phenomena for several statistics of conjugation-invariant permutations, with a dependence on the number of fixed points and sometimes $2$-cycles. Recent advances motivate this idea by generalizing asymptotic properties of uniformly random permutations to conjugation-invariant random permutations under certain hypotheses [Fér13, Kam18, FKL22, HR22]. The aim of this paper is to make further progress in this direction by introducing a “geometric” construction for random permutations in given conjugacy classes, see Section 1.2. This new point of view allows us to establish asymptotic results on the longest monotone subsequences, Robinson–Schensted shape, number of records, and pattern counts.
@@
The maximum length of a monotone subsequence in a uniform permutation famously behaves as $2\sqrt{n}$ [VK77]. Baik and Rains [BR01] showed similar asymptotics for random involutions, with a dependence on the number of fixed points, and Kammoun [Kam18] proved that the $2\sqrt{n}$ asymptotic holds for conjugation-invariant permutations with few cycles. In Section 2.1 we extend these results to general conjugation-invariant random permutations for their longest decreasing subsequences. Longest increasing subsequences are harder to handle, and we only provide such asymptotics up to a multiplicative constant.
@@
The Robinson–Schensted shape of a uniform permutation converges, after suitable rescaling, to a limit curve [LS77, VK77]. Kammoun [Kam18] extended this result to conjugation-invariant permutations with few cycles, and conjectured that this assumption could be lifted. In Section 2.2 we provide a positive answer to this conjecture.
@@
The numbers of records in uniform permutations are known to satisfy asymptotic normality. In Section 2.3 we prove a general limit theorem for the number of high records (or left-to-right maxima) in conjugation-invariant permutations, showcasing a phase transition as the number of fixed points grows. More precisely, the number of high records interpolates from logarithmic with Gaussian fluctuations to asymptotically Gamma, when the number of non-fixed points is of order $n/\sqrt{\log n}$. We also find the first order asymptotics of low records (or left-to-right minima), and their second order asymptotics in some cases.
@@
Janson et al. [JNZ15] established asymptotic normality for all pattern counts in uniform permutations. Kammoun [Kam22] and Hamaker and Rhoades [HR22] extended this to conjugation-invariant permutations, under certain assumptions on their cycles. The authors of [HR22] conjectured that the (non-degenerate) asymptotic normality of pattern counts holds for all conjugation-invariant permutations, where the asymptotic variance depends on the numbers of fixed points and $2$-cycles. In Section 2.4 we partially answer this conjecture, by proving asymptotic normality for all conjugation-invariant permutations, and non-degeneracy in some cases.
@@ 1.2 A geometric construction
Let $\mathcal{P}=\{Z_{1},\dots,Z_{n}\}$ be a finite subset of $\mathbb{R}^{2}$ with no redundant x- or y-coordinates. Then we can define a permutation $\tau$ of $[n]$ by letting $\tau(i)=j$ when the $i$-th point from the left in $\mathcal{P}$ is $j$-th from the bottom. We write $\tau={\rm Perm}\left(\mathcal{P}\right)$ for this permutation; see Figure 1 for an example.
Figure 1: Representation of the geometric construction of a uniform $(2,2,2,1)$-cyclic permutation as in 1.1, written in one-line notation and in cycle product notation.
By considering adequate planar point processes $\mathcal{P}$, we may obtain random permutations $\tau$ with various laws: for instance if $Z_{1},\dots,Z_{n}$ are i.i.d. ${\rm Unif}\left([0,1]^{2}\right)$ variables then $\tau$ is uniformly random, and if we symmetrize this family with respect to the diagonal of $[0,1]^{2}$ then $\tau$ is a uniform involution of size $2n$ with no fixed point [BR01]. An asset of this point of view is that some properties of $\tau$ may be easier to derive directly from $\mathcal{P}$: see e.g. [AD95] for a study of increasing subsequences in uniform permutations, and [Kiw06] for the case of involutions. 1.1 generalizes the geometric construction of Baik and Rains for random involutions [BR01] to uniform permutations in any conjugacy class. Although simple, this lemma is fundamental to our analysis: general conjugation–invariant permutations do not benefit from the same valuable algebraic properties as uniform permutations or involutions, and our geometric construction bypasses this problem.
If $t$ is a cycle type of size $n$, define $\mathcal{V}_{t}:=\left\{\big{.}(p,k,l)\,:\,p\in[n],k\in[t_{p}],l\in[p]\right\}$. Using a family of i.i.d. uniform variables indexed by this set, we can construct a uniform $t$-cyclic permutation as follows. This is illustrated on Figure 1.
@@
Let $t$ be a cycle type of size $n$. Let $\left(U_{p,k}^{l}\right)_{(p,k,l)\in\mathcal{V}_{t}}$ be a family of i.i.d. ${\rm Unif}\left([0,1]\right)$ random variables. For each $(p,k,l)\in\mathcal{V}_{t}$, write $Z_{p,k}^{l}:=\left(U_{p,k}^{l},U_{p,k}^{l+1}\right)$ where $l\!+\!1$ is taken modulo $p$, and set $\mathcal{P}:=\left\{Z_{p,k}^{l}:(p,k,l)\in\mathcal{V}_{t}\right\}$. Then $\tau:={\rm Perm}\left(\mathcal{P}\right)$ is a uniform $t$-cyclic permutation.
Proof.
The list of x-coordinates of $\mathcal{P}$ is the same as its list of y-coordinates, namely $\mathcal{U}=\left\{U_{p,k}^{l}\,:\,(p,k,l)\in\mathcal{V}_{t}\right\}$. Write $i_{p,k}^{l}\in[n]$ for the relative order of $U_{p,k}^{l}$ in $\cal U$, i.e. $U_{p,k}^{l}$ is the $i_{p,k}^{l}$-th lowest value in $\cal U$. Then the presence of $\left(U_{p,k}^{l},U_{p,k}^{l\!+\!1}\right)$ in $\mathcal{P}$ means that $\tau\left(i_{p,k}^{l}\right)=i_{p,k}^{l\!+\!1}$. Therefore $\tau$ is the product of cycles $\left(i_{p,k}^{1},\dots,i_{p,k}^{p}\right)$ for $p\in[n]$, $k\in[t_{p}]$. In other words, $\tau$ is obtained by writing the sequence $\left(i_{p,k}^{l}\right)_{(p,k,l)\in\mathcal{V}_{t}}$ with the lexicographic order on $(p,k,l)$, parenthesizing it according to the cycle type $t$, and interpreting this as a product of cycles. By standardization of the i.i.d. sequence $\left(U_{p,k}^{l}\right)_{(p,k,l)\in\mathcal{V}_{t}}$, the sequence $\left(i_{p,k}^{l}\right)_{(p,k,l)\in\mathcal{V}_{t}}$ is the one-line representation of a uniformly random permutation of $[n]$. From there, it is classical that $\tau$ is a uniform $t$-cyclic permutation. ∎
@@
As we explain in Sections 2.1, 2.2 and 2.3, monotone subsequences and records of $\tau={\rm Perm}\left(\mathcal{P}\right)$ can be read directly on $\mathcal{P}$. This allows us to study these statistics of the permutation by investigating the behavior of the point process in certain regions of the plane. To this aim, a list of useful properties of the geometric construction may be found in Section 3. Regarding pattern counts, we rather use the “weak dependency” of points in $\mathcal{P}$ (see Section 2.4 for more details).
@@ 1.3 Notation
We denote by $\mathfrak{S}_{n}$ the set of permutations of $[n]$ and by $\Delta$ the diagonal of $[0,1]^{2}$. Throughout this paper we consider the standard partial order on $\mathbb{R}^{2}$ defined by $(x,y)\leq(x^{\prime},y^{\prime})$ when $x\leq x^{\prime}$ and $y\leq y^{\prime}$.
Let $t$ be a cycle type of size $n$. We say that $\mathcal{P}$ is a geometric construction of $t$ when $\mathcal{P}$ is constructed as in 1.1. We can decompose it into its cycles $\mathcal{P}=\bigcup_{p\in[n],k\in[t_{p}]}\mathcal{P}_{p,k}$ where $\mathcal{P}_{p,k}:=\left\{Z_{p,k}^{l}:l\in[p]\right\}$. We can also decompose it into $\mathcal{P}_{\Delta}:=\mathcal{P}\cap\Delta$ and $\check{\mathcal{P}}:=\mathcal{P}\setminus\Delta$, and we write $\check{t}:=(0,t_{2},t_{3},\dots,t_{n})$ for the cycle type of size $\check{n}:=n-t_{1}$ with no fixed point. It is straightforward to check that $\check{\mathcal{P}}$ is a geometric construction of $\check{t}$.
We define a “successor” operator $\mathfrak{s}:(p,k,l)\in\mathcal{V}_{t}\mapsto(p,k,l\!+\!1)\in\mathcal{V}_{t}$ where $l\!+\!1$ is taken modulo $p$, and a graph $\mathcal{L}_{t}:=(\mathcal{V}_{t},\mathcal{E}_{t})$ with vertice $\mathcal{V}_{t}$ and edges between $i$ and $\mathfrak{s}(i)$ for each $i\in\mathcal{V}_{t}$ (forgetting self-loops). We refer to $\mathcal{L}_{t}$ as the “dependency graph” associated with $t$. See Figure 2 for an example, and Section 7.1 for an explanation of this name.
@@ 2.1 Longest monotone subsequences
The study of monotone subsequences in permutations has been an active and captivating field of research; see [Rom15] for an introduction. In particular when $\tau_{n}$ is uniformly random, the law of $\mathrm{LIS}(\tau_{n})$ (which is the same as that of $\mathrm{LDS}(\tau_{n})$ in this case) is now well understood: it behaves asymptotically as $2\sqrt{n}$ [VK77], satisfies a large deviation principle [DZ99], and admits Tracy–Widom fluctuations of order $n^{1/6}$ [BDJ99].
One might expect these results to mostly hold when the law of $\tau_{n}$ is invariant by conjugation. The first advance on this question was obtained by Baik and Rains, who studied random involutions [BR01]. More recently, Kammoun [Kam18, Kam23] and Guionnet [GK23] could settle the case of conjugation-invariant random permutations with few cycles. Here we start by stating first order asymptotic results and concentration inequalities for the longest decreasing subsequences of uniform permutations with given cycle types and no fixed points.
@@
$\left\{\begin{array}[]{ll}\mathbf{P}\left(\mathrm{LDS}(\tau_{n})<2(1-\delta)\sqrt{n-t_{1}^{(n)}}\right)\leq\mathbb{E}\left[\exp\left(\big{.}-c_{\delta}\left(n-t_{1}^{(n)}\right)\right)\right];\vspace{.5em}\\ \mathbf{P}\left(\mathrm{LDS}(\tau_{n})>2(1+\delta)\sqrt{n-t_{1}^{(n)}}\right)\leq\mathbb{E}\left[\exp\left(-c_{\delta}\sqrt{n-t_{1}^{(n)}}\right)\right].\end{array}\right.$
In particular if $n-t_{1}^{(n)}\to\infty$ in probability then $\frac{1}{\sqrt{n-t_{1}^{(n)}}}\mathrm{LDS}(\tau_{n})\to 2$ in probability and in $L^{q}$ for all $q\geq 1$.
Remark 1.
In comparison with [Kam18, Theorem 1.2] and its variants in [Kam23], 2.2 does not require any hypothesis on the number of cycles. Our results can in particular be applied to random permutations with cycle weights, some of which have a lot of cycles [BUV11, EU14], and to central virtual permutations, some of which have a macroscopic number of fixed points [Tsi00]. A similar observation can be made for all of the results presented in this paper.
@@
$\left\{\begin{array}[]{ll}\mathbf{P}\left(\big{.}\mathrm{LIS}(\tau_{n})<2(1-\delta)\sqrt{n}\right)\leq\exp\left(-c_{\delta}n\right);\vspace{.5em}\\ \mathbf{P}\left(\mathrm{LIS}(\tau_{n})>t_{1}^{(n)}+2(1+\delta)\sqrt{3\left(n-t_{1}^{(n)}\right)}\right)\leq\mathbb{E}\left[\exp\left(-c_{\delta}\sqrt{n-t_{1}^{(n)}}\right)\right].\end{array}\right.$
@@
Let us briefly explain the techniques used to prove the results of this section. Deuschel and Zeitouni [DZ95] studied the first order asymptotics of longest monotone subsequences in ${\rm Perm}\left(Z_{1},\dots,Z_{n}\right)$, where $Z_{1},\dots,Z_{n}$ are i.i.d. under certain probability densities on $[0,1]^{2}$ (such random permutations are called locally uniform in [Sjö23]). This is made possible by the following key observation: increasing and decreasing subsequences of $\tau={\rm Perm}\left(\mathcal{P}\right)$ correspond to up-right and down-right paths of points in $\mathcal{P}$. The authors of [DZ95] then slice the unit square into small rectangles in which the points are almost uniformly distributed, so that they may locally apply the result of [VK77], and finally stitch these local paths of points into one global path. When studying uniform permutations in given conjugacy classes, the same method can be adapted to the geometric construction of 1.1.
Finally, we state a reasonable conjecture in the direction of 2.3 and 2.4. It was already established for random involutions in [BR01], and for conjugation-invariant random permutations satisfying certain cycle constraints in [Kam18, Theorem 1.2] and [Kam23, Theorem 3].
@@
Baik and Rains actually proved [BR01, Theorems 3.2 and 3.4] that if $\tau_{n}$ is a uniform $t^{(n)}$-cyclic permutation where $t^{(n)}~{}=~{}\left(t_{1}^{(n)},t_{2}^{(n)},0,\dots\right)$ satisfies $t_{1}^{(n)}/\sqrt{n}\to\alpha\geq 0$, i.e. if $\tau_{n}$ is a uniform involution with $\alpha\sqrt{n}+o(\sqrt{n})$ fixed points, then
$\frac{1}{\sqrt{n}}\mathrm{LIS}(\tau_{n})\underset{n\to\infty}{\longrightarrow}\left\{\begin{array}[]{ll}2&\text{if }\alpha\leq 1\vspace{.5em}\\ \alpha+\frac{1}{\alpha}&\text{if }\alpha\geq 1\end{array}\right.$
in probability. We could expect the same first order asymptotics for any sequence of conjugation-invariant random permutations such that $t_{1}^{(n)}/\sqrt{n}\to\alpha$ in probability. However we chose not to include this refinement of 2.5 in its statement, as it might be more speculative.
The results of this section are proved in Section 4.

### v4

@@ Abstract
Acknowledgments: The author is very grateful to Valentin Féray for insightful discussions and valuable suggestions, and to Mohamed Slim Kammoun for helpful comments.
@@ 1.1 Conjugation-invariant random permutations
We say that a random permutation $\tau$ of $n]:=\{1,\dots,n\}$ is conjugation-invariant when for any fixed permutation $\rho$, the conjugate $\rho\circ\tau\circ\rho^{-1}$ follows the same law as $\tau$. Standard examples of conjugation-invariant permutations are uniformly random permutations, random involutions, Ewens random permutations [E72], their generalizations [BUV11, K78, T00], and many more.
@@
There are reasons to believe in universality phenomena for several statistics of conjugation-invariant permutations, with a dependence on the number of fixed points and sometimes $2$-cycles. Recent advances motivate this idea by generalizing asymptotic properties of uniformly random permutations to conjugation-invariant random permutations under certain hypotheses [F13, FKL22, HR22, K18]. The aim of this paper is to make further progress in this direction by introducing a “geometric” construction for random permutations in given conjugacy classes, see ˜1.3 in Section˜1.2. This new point of view allows us to establish asymptotic results on the longest monotone subsequences, Robinson–Schensted shape, number of records, and pattern counts.
@@
The maximum length of a monotone subsequence in a uniform permutation famously behaves as $2\sqrt{n}$ [VK77]. Baik and Rains [BR01] showed similar asymptotics for random involutions, with a dependence on the number of fixed points, and Kammoun [K18] proved that the $2\sqrt{n}$ asymptotic holds for conjugation-invariant random permutations with few cycles. In Section˜2.1 we extend these results to general conjugation-invariant permutations for their longest decreasing subsequences. Longest increasing subsequences are harder to handle, and we only provide such asymptotics up to a multiplicative constant.
@@
The Robinson–Schensted shape of a uniform permutation converges, after suitable rescaling, to a limit curve [LS77, VK77]. Kammoun [K18] extended this result to conjugation-invariant permutations with few cycles, and conjectured that this assumption could be lifted. In Section˜2.2 we provide a positive answer to this conjecture.
@@
The numbers of records in uniform permutations are known to satisfy asymptotic normality. In Section˜2.3 we prove a general limit theorem for the number of high records (or left-to-right maxima) in conjugation-invariant permutations, showcasing a phase transition as the number of fixed points grows. More precisely, the number of high records interpolates from logarithmic with Gaussian fluctuations to asymptotically Gamma, when the number of non-fixed points is of order $n/\sqrt{\log n}$. We also find the first order asymptotics of low records (or left-to-right minima), and their second order asymptotics in some cases.
@@
Janson et al. [JNZ15] established asymptotic normality for all pattern counts in uniform permutations. Kammoun [K22] and Hamaker and Rhoades [HR22] extended this to conjugation-invariant permutations, under certain assumptions on their cycles. The authors of [HR22] conjectured that the (non-degenerate) asymptotic normality of pattern counts holds for all conjugation-invariant permutations, where the asymptotic variance depends on the numbers of fixed points and $2$-cycles. In Section˜2.4 we partially answer this conjecture, by proving asymptotic normality for all conjugation-invariant permutations, and non-degeneracy in some cases.
Our results can be applied to most models of conjugation-invariant permutations, with virtually no restrictions on their cycle types. For example, they are novel for random permutations with cycle weights, some of which have a lot of cycles [BUV11, EU14], and for central virtual permutations, some of which have a macroscopic number of fixed points [T00].
@@ 1.2 A geometric construction
We begin by recalling the notion of standardization: if $(y_{1},\dots,y_{n})$ is a sequence of pairwise distinct numbers, define $\tau=\mathrm{st}\langle y_{1},\dots,y_{n}\rangle$ as the unique permutation of $[n]$ such that:
$\text{for all }1\leq i,j\leq n,\quad\tau(i)<\tau(j)\Leftrightarrow y_{i}<y_{j}.$
The following two results are classical and easy to check:
@@
Let $\mu$ be an atomless probability distribution on $\mathbb{R}$, and let $Y_{1},\dots,Y_{n}$ be i.i.d. random variables distributed under $\mu$. Then $Y_{1},\dots,Y_{n}$ are a.s. pairwise distinct, and the permutation $\mathrm{st}\langle Y_{1},\dots,Y_{n}\rangle$ is uniformly random in $\mathfrak{S}_{n}$.
Lemma 1.2.
Let $t$ be a cycle type of size $n$, and let $\mathfrak{s}$ be an arbitrary fixed $t$-cyclic permutation. If $\sigma$ is a uniformly random permutation of $\mathfrak{S}_{n}$ then $\sigma\circ\mathfrak{s}\circ\sigma^{-1}$ is a uniformly random $t$-cyclic permutation.
Now, consider a finite subset $\mathcal{P}=\{Z_{1}=(X_{1},Y_{1}),\dots,Z_{n}=(X_{n},Y_{n})\}$ of $\mathbb{R}^{2}$ with no redundant x- or y-coordinates. We can define a permutation $\tau$ of $[n]$ by letting $\tau(i)=j$ when the $i$-th point from the left in $\mathcal{P}$ is the $j$-th point from the bottom; equivalently, $\tau=\mathrm{st}\langle Y_{(1)},\dots,Y_{(n)}\rangle$ where $Z_{(1)},\dots,Z_{(n)}$ are ordered such that $X_{(1)}<\dots<X_{(n)}$. We denote by $\tau={\rm Perm}\left(\mathcal{P}\right)$ this permutation.
By considering adequate planar point processes $\mathcal{P}$, we may obtain random permutations $\tau$ with various laws: for instance if $Z_{1},\dots,Z_{n}$ are i.i.d. ${\rm Unif}\left([0,1]^{2}\right)$ variables then $\tau$ is uniformly random, and if we symmetrize this family with respect to the diagonal of $0,1]^{2}$ then $\tau$ is a uniform involution of size $2n$ with no fixed point. The latter construction was notably used by Baik and Rains in [BR01]. An asset of this point of view is that some properties of $\tau$ may be easier to derive directly from $\mathcal{P}$: see e.g. [AD95] for a study of increasing subsequences in uniform permutations, and [K06] for the case of involutions. ˜1.3 generalizes the geometric construction of Baik and Rains for random involutions [BR01] to uniform permutations in any conjugacy class. Although simple, this lemma is fundamental to our analysis: general conjugation–invariant permutations do not benefit from the same valuable algebraic properties as uniform permutations or involutions, and our geometric construction bypasses this problem.
Lemma 1.3.
@@
Let $t$ be a cycle type of size $n$, and let $\mathfrak{s}\in\mathfrak{S}_{n}$ be an arbitrary $t$-cyclic permutation. Let $\left(U_{i}\right)_{i\in[n]}$ be a sequence of i.i.d. ${\rm Unif}\left([0,1]\right)$ random variables. Set $Z_{i}:=\left(U_{i},U_{\mathfrak{s}(i)}\right)$ for each $i\in[n]$, and $\mathcal{P}:=\left\{Z_{i},i\in[n]\right\}$. Then $\tau:={\rm Perm}\left(\mathcal{P}\right)$ is a uniform $t$-cyclic permutation.
Figure 1: The geometric construction of a random $t$-cyclic permutation for $t=(1,2,1,1)$. Here we used the arbitrary $t$-cyclic permutation $\mathfrak{s}:=(1)\circ(2,3)\circ(4,5)\circ(6,7,8)\circ(9,10,11,12)$. The resulting permutation $\tau$ is written in one-line notation and in cycle product notation, and its cycles are highlighted on the point set. For example, the points $(U_{6},U_{7}),(U_{7},U_{8}),(U_{8},U_{6})$ are linked by the red dotted line, and correspond to the 33-cycle $(9,4,12)$ in $\tau$.
See Figure˜1 for an illustration. In practice, instead of an arbitrary $t$-cyclic permutation $\mathfrak{s}$, we will use a canonical one on an ad hoc set of indices: let $\mathcal{V}_{t}:=\left\{\big.(p,k,l)\,:\,p\in[n],k\in[t_{p}],l\in[p]\right\}$, and $\mathfrak{s}:(p,k,l)\mapsto(p,k,l\!+\!1)$ where l+1l\!+\!1 is taken modulo $p$. Then the construction of ˜1.3 works with a family $\left(U_{p,k}^{l}\right)_{(p,k,l)\in\mathcal{V}_{t}}$ of i.i.d. ${\rm Unif}\left([0,1]\right)$ random variables, the points $Z_{p,k}^{l}:=\left(U_{p,k}^{l},U_{p,k}^{l+1}\right)$, and the point set $\mathcal{P}:=\left\{Z_{p,k}^{l}:(p,k,l)\in\mathcal{V}_{t}\right\}$. With this notation, we say that $\mathcal{P}$ is a geometric construction of $t$.
Proof of ˜1.3.
Let $\sigma=(\sigma_{1},\dots,\sigma_{n}):=\mathrm{st}\langle U_{1},\dots,U_{n}\rangle$, i.e. $U_{i}$ is the $\sigma_{i}$-th lowest number in $\mathcal{U}:=\left\{U_{i},i\in[n]\right\}$. Observe that $\mathcal{U}$ is the list of x-coordinates of $\mathcal{P}$ and its list of y-coordinates as well; consequently, $U_{i}$ is the $\sigma_{i}$-th lowest x-coordinate of points in $\mathcal{P}$ and $U_{\mathfrak{s}(i)}$ is the $\sigma_{\mathfrak{s}(i)}$-th lowest y-coordinate of points in $\mathcal{P}$.. Since $\left(U_{i},U_{\mathfrak{s}(i)}\right)$ is in $\mathcal{P}$, this precisely means that $\tau(\sigma_{i})=\sigma_{\mathfrak{s}(i)}$. Therefore, $\tau=\sigma\circ\mathfrak{s}\circ\sigma^{-1}$. Using Lemmas 1.1 and 1.2, this concludes the proof. ∎
@@
As we explain in Sections˜2.1, 2.2 and 2.3, monotone subsequences and records of $\tau={\rm Perm}\left(\mathcal{P}\right)$ can be read directly on $\mathcal{P}$. This allows us to study these statistics of the permutation by investigating the behavior of the point process in certain regions of the plane. To this aim, a list of useful properties of the geometric construction may be found in Section˜3. Regarding pattern counts, we rather use the “weak dependency” of points in $\mathcal{P}$ (see Section˜2.4 for more details).
@@ 1.3 Notation
We denote by $\mathfrak{S}_{n}$ the set of permutations of $[n]$ and by $\Delta$ the diagonal of $0,1]^{2}$. Throughout this paper we consider the standard partial order on $\mathbb{R}^{2}$ defined by $(x,y)\leq(x^{\prime},y^{\prime})$ when $x\leq x^{\prime}$ and $y\leq y^{\prime}$.
Let $t$ be a cycle type of size $n$, and let $\mathcal{P}$ be a geometric construction of $t$. We can decompose it into its cycles $\mathcal{P}=\bigcup_{p\in[n],k\in[t_{p}]}\mathcal{P}_{p,k}$ where $\mathcal{P}_{p,k}:=\left\{Z_{p,k}^{l}:l\in[p]\right\}$. We can also decompose it into $\mathcal{P}_{\Delta}:=\mathcal{P}\cap\Delta$ and $\check{\mathcal{P}}:=\mathcal{P}\setminus\Delta$, and we write $\check{t}:=(0,t_{2},t_{3},\dots,t_{n})$ for the cycle type of size $\check{n}:=n-t_{1}$ with no fixed point. It is straightforward to check that $\check{\mathcal{P}}$ is a geometric construction of $\check{t}$.
Define a graph $\mathcal{L}_{t}:=(\mathcal{V}_{t},\mathcal{E}_{t})$ with vertice $\mathcal{V}_{t}$ and edges between $i$ and $\mathfrak{s}(i)$ for each $i\in\mathcal{V}_{t}$ (forgetting self-loops). We refer to $\mathcal{L}_{t}$ as the “dependency graph” associated with $t$. See Figure˜2 for an example, and Section˜7.1 for an explanation of this name.
@@ 2.1 Longest monotone subsequences
The study of monotone subsequences in permutations has been an active and captivating field of research; see [R15] for an introduction. In particular when $\tau_{n}$ is uniformly random, the law of $\mathrm{LIS}(\tau_{n})$ (which is the same as that of $\mathrm{LDS}(\tau_{n})$ in this case) is now well-understood: it behaves asymptotically as $2\sqrt{n}$ [VK77], satisfies a large deviation principle [DZ99, S98], and admits Tracy–Widom fluctuations of order $n^{1/6}$ [BDJ99].
One might expect these results to mostly hold when the law of $\tau_{n}$ is conjugation-invariant. The first advance on this question was obtained by Baik and Rains, who studied random involutions [BR01]. More recently, Kammoun [K18, K23] and Guionnet [GK23] could settle the case of conjugation-invariant random permutations with few cycles. Here we start by stating first order asymptotic results and concentration inequalities for the longest decreasing subsequences of uniform permutations with given cycle types and no fixed points.
@@
$\left\{\begin{array}[]{ll}\mathbf{P}\left(\mathrm{LDS}(\tau_{n})<2(1-\delta)\sqrt{n-t_{1}^{(n)}}\right)\leq\mathbb{E}\left[\exp\left(\big.-c_{\delta}\left(n-t_{1}^{(n)}\right)\right)\right];\vskip 5.0pt\\ \mathbf{P}\left(\mathrm{LDS}(\tau_{n})>2(1+\delta)\sqrt{n-t_{1}^{(n)}}\right)\leq\mathbb{E}\left[\exp\left(-c_{\delta}\sqrt{n-t_{1}^{(n)}}\right)\right].\end{array}\right$.
In particular if $n-t_{1}^{(n)}\to\infty$ in probability then $\frac{1}{1+\sqrt{n-t_{1}^{(n)}}}\mathrm{LDS}(\tau_{n})\to 2$ in probability and in $L^{q}$ for all $q\geq 1$.
@@
$\left\{\begin{array}[]{ll}\mathbf{P}\left(\big.\mathrm{LIS}(\tau_{n})<2(1-\delta)\sqrt{n}\right)\leq\exp\left(-c_{\delta}n\right);\vskip 5.0pt\\ \mathbf{P}\left(\mathrm{LIS}(\tau_{n})>t_{1}^{(n)}+2(1+\delta)\sqrt{3\left(n-t_{1}^{(n)}\right)}\right)\leq\mathbb{E}\left[\exp\left(-c_{\delta}\sqrt{n-t_{1}^{(n)}}\right)\right].\end{array}\right$.
@@
Let us briefly explain the techniques used to prove the results of this section. Deuschel and Zeitouni [DZ95] studied the first order asymptotics of longest monotone subsequences in ${\rm Perm}\left(Z_{1},\dots,Z_{n}\right)$, where $Z_{1},\dots,Z_{n}$ are i.i.d. under certain probability densities on $0,1]^{2}$ (such random permutations are called locally uniform in [S23]). This is made possible by the following key observation: increasing and decreasing subsequences of $\tau={\rm Perm}\left(\mathcal{P}\right)$ correspond to up-right and down-right paths of points in $\mathcal{P}$. The authors of [DZ95] then slice the unit square into small rectangles in which the points are almost uniformly distributed, so that they may locally apply the result of [VK77], and finally stitch these local paths of points into one global path. When studying uniform permutations in given conjugacy classes, the same method can be adapted to the geometric construction of ˜1.3.
Finally, we state a reasonable conjecture in the direction of ˜2.3 and ˜2.4. It was already established for random involutions in [BR01], and for conjugation-invariant random permutations satisfying certain cycle constraints in [K18, Theorem 1.2] and [K23, Theorem 3].
@@
Baik and Rains actually proved [BR01, Theorems 3.2 and 3.4] that if $\tau_{n}$ is a uniform $t^{(n)}$-cyclic permutation where $t^{(n)}~\hskip-3.00003pt=~\hskip-3.00003pt\left(t_{1}^{(n)},t_{2}^{(n)},0,\dots\right)$ satisfies $t_{1}^{(n)}/\sqrt{n}\to\alpha\geq 0$, i.e. if $\tau_{n}$ is a uniform involution with $\alpha\sqrt{n}+o(\sqrt{n})$ fixed points, then
$\frac{1}{\sqrt{n}}\mathrm{LIS}(\tau_{n})\underset{n\to\infty}{\longrightarrow}\left\{\begin{array}[]{ll}2&\text{if }\alpha\leq 1\vskip 5.0pt\\ \alpha+\frac{1}{\alpha}&\text{if }\alpha\geq 1\end{array}\right$.
in probability. We could expect the same first order asymptotics for any sequence of conjugation-invariant random permutations such that $t_{1}^{(n)}/\sqrt{n}\to\alpha$ in probability. However we chose not to include this refinement of ˜2.5 in its statement, as it might be more speculative.
The results of this section are proved in Section˜4.